complete the pattern 5,9,36,246,3036,?.

Grade:10

1 Answers

Saurabh Koranglekar
5 years ago
Dear student

Please consider the approach

(5+1)*1.5=9 ; 1.5 → 3 → 6....

(9+3)*3=36; 1 → 3 → 5 → 7........

(36+5)*6=246

(246+7)*12=3036

(3036+9)*24=73080

Hope it helps

Regards